OpenStudy (anonymous):

Greg Lawrence anticipates he will need approximately $218,700 in 10 years to cover his 3-year-old daughter’s college bills for a 4-year degree. How much would he have to invest today, at an interest rate of 7 percent compounded semiannually? (Please use only the following provided Table) (Round your answer to 2 decimal places. Omit the "$" sign in your response.)

OpenStudy (anonymous):

http://www.purplemath.com/modules/expofcns4.htm \[218,700=P \left(1+ \frac{ .07 }{ 2 } \right)^{2*10}\]

OpenStudy (anonymous):

Basically, this is the formula you looking at. You need to solve for P.

OpenStudy (anonymous):

So, you'll want to isolate P. \[\frac{ 218,700 }{ \left( 1+\frac{ .07 }{ 2 } \right)^{20}}=P\]

OpenStudy (anonymous):

So if my math is right, you'll have to invest $109911.16